Provincial​ Imports, Inc., has assembled past ​(2019​) financial statements​ (income statement and balance sheet LOADING...​) and financial projections for use in preparing financial plans for the coming year ​(2020​). Information related to financial projections for the year 2020 is as​ follows: ​

(1) Projected sales are $ 5,992,000. ​

(2) Cost of goods sold in 2019 includes $ 1,000,000 in fixed costs.

​(3) Operating expense in 2019 includes $ 245,000 in fixed costs. ​

(4) Interest expense will remain unchanged. ​

(5) The firm will pay cash dividends amounting to 35 % of net profits after taxes. ​

(6) Cash and inventories will double. ​

(7) Marketable​ securities, notes​ payable, long-term​ debt, and common stock will remain unchanged. ​

(8) Accounts​ receivable, accounts​ payable, and other current liabilities will change in direct response to the change in sales.

​(9) A new computer system costing $ 364,000 will be purchased during the year. Total depreciation expense for the year will be $ 118,000. ​

(10) The tax rate will remain at 21 %.

a. Prepare a pro forma income statement for the year ended December​ 31, 2020​, using the fixed cost data given to improve the accuracy of the​ percent-of-sales method.

b. Prepare a pro forma balance sheet as of December​ 31, 2020​, using the information given and the judgmental approach. Include a reconciliation of the retained earnings account.

c. Analyze these​ statements, and discuss the resulting external financing required.

Provincial​ Imports, Inc. Income Statement for the Year Ended December​ 31, 2019

Sales revenue $ 5,007,000 ​

Less: Cost of goods sold 2,741,000

Gross profits $ 2,266,000 ​

Less: Operating expenses 854,000

Operating profits $ 1,412,000

​Less: Interest expense 206,000

Net profits before taxes $ 1,206,000

​Less: Taxes ​(rate equals 21 %​) 253,260

Net profits after taxes $ 952,740 ​

Less: Cash dividends 333,459

To retained earnings $ 619,281

Provincial​ Imports, Inc. Balance Sheet December​ 31, 2019

Assets Liabilities and​ Stockholders' Equity Cash $ 204,000

Accounts payable $ 700,000

Marketable securities 225,000

Taxes payable 95,000

Accounts receivable 618,000

Notes payable 192,000

Inventories 495,000

Other current liabilities 4,800

Total current assets $ 1,542,000

Total current liabilities $ 991,800

Net fixed assets 1,406,000 ​

Long-term debt 498,200

Common stock 74,000

Retained earnings 1,384,000

Total assets $ 2,948,000

Total liabilities and equity $ 2,948,000

a. Prepare a pro forma income statement for the year ended December​ 31, 2020​, using the fixed cost data given to improve the accuracy of the​ percent-of-sales method. Complete the pro forma income statement for the year ended December​ 31, 2020 ​below:  ​(Round to the nearest​ dollar.) Pro Forma Income Statement Provincial Imports, Inc. for the Year Ended December 31, 2020 (percent-of-sales method)

Sales:

Less: Cost of goods sold:

Gross Profits:

Less: Operating Expenses:

Operating Profits:

Less: Interest Expense:

Net profits before taxes:

Less: Taxes (rate = 21%):

Net profits after taxes:

Less: Cash dividends (35%):

To Retained Earnings:

b. Prepare a pro forma balance sheet as of December​ 31, 2020​, using the information given and the judgmental approach. Include a reconciliation of the retained earnings account. ​ NOTE: Taxes payable for 2019 are about 37.5109 % of the  2019taxes on the income statement. The pro forma value is obtained by taking 37.5109 % of the 2020 taxes. Complete the assets part of the pro forma balance sheet as of December​ 31, 2020​: ​(Round to the nearest​ dollar.) Pro Forma Balance Sheet Provincial Imports, Inc. for the Year Ended December 31, 2020 (Judgmental Method)

Cash:

Marketable securities:

Accounts receivable:

Inventories:

Total current assets:

Net fixed assets:

Total assets:

Complete the liabilities and equity part of the pro forma balance sheet as of December​ 31, 2020​: ​(Round to the nearest​ dollar.) Pro Forma Balance Sheet Provincial Imports, Inc. for the Year Ended December 31, 2020 (Judgmental Method)

Accounts payable:

Taxes payable:

Notes payable:

Other current liabilities:

Total current liabilities:

Long-term debt:

Common stock:

Retained earnings:

External funds required:

Total liabilities and stockholders' equity:

c. Using the judgmental​ approach, the external funds requirement is ​$
Answer #1

Part A

Pro Forma Income Statement

Provincial Imports, Inc.

for the Year Ended December 31, 2020

(percent-of-sales method)

Sales

5992000

Less: Cost of goods sold (34.7713%*5992000)+ $1,000,000)

3083496

Gross profits

2908504

Less: Operating expenses (12.1630%*5992000 + $245,000)

973807

Operating profits

1934697

Less: Interest expense

206000

Net profits before taxes

1728697

Less: Taxes (21% ´ NPBT)

363026

Net profits after taxes

1365671

Less: Cash dividends (35% ´ NPAT)

477985

To Retained earnings

887686

Cost of goods sold rate = 1741000/5007000 = 34.7713%

Operating expenses rate = (854000-245000)/5007000 = 12.16% = 12%

Part B

Pro Forma Balance Sheet

Provincial Imports, Inc.

December 31, 2013

(Judgmental Method)

Assets

Liabilities and Equity

Cash

408000

Accounts payable

837708

Marketable securities

225000

Taxes payable

136174

Accounts receivable

739576

Notes payable

192000

Inventories

990000

Other current liabilities

5744

Current assets

2362576

Current liabilities

1171626

Net fixed assets

1652000

Long-term debt

498200

Common stock

74000

Retained earnings

2271686

External funds required

(936)

Total assets

4014576

Total liabilities and stockholders’ equity

4014576

618000*119.6725% = 739576

1406000+364000-118000 = 1652000

700000*119.6725% = 837708

4800*5992000/5007000 = 5744

363026*37.5109% = 136174

Retained earnings = 1384000+887686 = 2271686

Part C

Using the judgmental approach, the external funds requirement is $(936)